Monthly Payments and Finance Charges. Kimberly Jensen of Storm Lake, Iowa, wants to buy some living room furniture for her new apartment. A local store offered credit at an APR of 16 percent, with a maximum term of four years. The furniture she wishes to purchase costs $4,800, with no down payment required. Make the following calculations:

a) What is the amount of the monthly payment if she borrowed for four years?

b) What are the total finance charges over that four year period?

Hint: (PMT x 48)

c) How would the payment change if she could afford a down payment of $500 with four years of financing?

d) What are the total finance charges over that four year period given the $500 down payment?
Answer #1

a)

What is the amount of the monthly payment

=4800/((1-(1+(16%/12))^(-4*12))/(16%/12))

=136.03

b)

What are the total finance charges

=136.03*48-4800

=1729.44

c)

new monthly payment is

=(4800-500)/((1-(1+(16%/12))^(-4*12))/(16%/12))

=121.86

d)

What are the total finance charges

=121.86*48-4300

=1549.28

the above is answer..